### v2

# From pre-transit to post-eclipse: investigating the impact of 3D temperature, chemistry, and dynamics on high-resolution emission spectra of the ultra-hot Jupiter WASP-76b

Abstract
High-resolution spectroscopy has provided a wealth of information about the climate and composition of ultra-hot Jupiters. However, the 3D structure of their atmospheres makes observations more challenging to interpret, necessitating 3D forward-modeling studies. In this work, we model phase-dependent thermal emission spectra of the archetype ultra-hot Jupiter WASP-76b to understand how the line strengths and Doppler shifts of Fe, CO, H2O, and OH evolve throughout the orbit. We post-process outputs of the SPARC/MITgcm global circulation model with the 3D Monte-Carlo radiative transfer code gCMCRT to simulate emission spectra at 36 orbital phases. We then cross-correlate the spectra with different templates to obtain CCF and – maps. For each species, our models produce consistently negative offsets in pre- and post-eclipse, which are driven by planet rotation. The size of these offsets is similar to the equatorial rotation velocity of the planet. Furthermore, we demonstrate how the weak vertical temperature gradient on the nightside of ultra-hot Jupiters mutes the absorption features of CO and H2O, which significantly hampers their detectability in pre- and post-transit. We also show that the and offsets in pre- and post-transit are not always a measure for the line-of-sight velocities in the atmosphere. This is because the cross-correlation signal is a blend of dayside emission and nightside absorption features. Finally, we highlight that the observational uncertainty in the known orbital velocity of ultra-hot Jupiters can be multiple km/s, which makes it hard for certain targets to meaningfully report absolute offsets.
1 Introduction
Owing to their short orbital periods, high equilibrium temperatures, and puffy atmospheres, ultra-hot Jupiters (UHJs; Parmentier et al. 2018; Bell & Cowan 2018; Arcangeli et al. 2018) are prime targets for atmospheric characterization via ground-based high-resolution spectroscopy. With the arrival of a host of new instruments including ESPRESSO (Pepe et al., 2021), IGRINS (Mace et al., 2018), IGRINS-2 (Oh et al., 2024), MAROON-X (Seifahrt et al., 2018), NIRPS (Artigau et al., 2024), CRIRES+ (Dorn et al., 2023), KPIC (Delorme et al., 2021), and GHOST (Kalari et al., 2024), we are now able to study the climate, chemical composition, and formation history of UHJs in unprecedented detail. Yet, one aspect that complicates the interpretation of the spectra of these planets is the inherent 3D structure of their atmospheres.
UHJs are tidally locked, such that they have a permanent dayside ( 3000 K) and a permanent nightside ( 1000 K). The dayside temperature is so high that molecules like H2 and H2O are subject to thermal dissociation, resulting in significant amounts of H, H-, and OH (Parmentier et al. 2018; Yan et al. 2021; Nugroho et al. 2021). Furthermore, a large fraction of metals become ionized (Hoeijmakers et al. 2019; Merritt et al. 2021; Prinoth et al. 2024). The absorption of intense stellar radiation by metal species, and potentially by TiO and VO, also causes the dayside temperature profile to be strongly inverted, such that temperature increases with altitude (Fortney et al. 2008; Haynes et al. 2015; Pino et al. 2020). On the dark nightside, however, the temperature profile should be non-inverted. Also, the lower nightside temperature allows for H2 and H2O to persist, while causing refractory species to condense into clouds (Helling et al. 2021; Komacek et al. 2022).
When it comes to the signature of “3D-ness” in the transmission spectra of UHJs, a relatively coherent picture has emerged. Absorption lines of species that probe the dayside (e.g., metals and CO) tend to show an increasing blueshift during the transit (Ehrenreich et al. 2020; Kesseli et al. 2022; Savel et al. 2023; Wardenier et al. 2023; Pelletier et al. 2023; Prinoth et al. 2023; Beltz et al. 2023; Simonnin et al. 2024; Seidel et al. 2025). This is a consequence of the strong day-night contrast of the planet in combination with tidally locked rotation. The day-to-night flow produces an additional negative velocity offset of a few km/s. As a result, the detection peaks of these “dayside” species occur at smaller and values compared to the expected planet position in the – map of the observation (Wardenier et al. 2023). For species that mainly probe the nightside of the planet (e.g., H2O), the Doppler shifts of the absorption lines can either become more negative or more positive during the transit (Wardenier et al. 2023, 2024), depending on the atmospheric drag strength. On top of this behavior, asymmetries in scale height (i.e., temperature), cloud cover, or the degree of ionization between the planet’s morning and evening terminator can cause the absorption line strengths to vary substantially during the transit (Wardenier et al. 2021; Savel et al. 2022; Gandhi et al. 2022; Prinoth et al. 2023). Typically, UHJs have a hotter morning limb due to an eastward circulation pattern.
With regard to emission spectroscopy, a number of studies have simulated high-resolution thermal emission spectra of hot and ultra-hot Jupiters based on 3D global circulation models (GCMs) (Zhang et al. 2017; Harada et al. 2021; Malsky et al. 2021; Beltz et al. 2021, 2022; Lee et al. 2022a; Van Sluijs et al. 2023; Beltz & Rauscher 2024). The common denominator between many of these works is the finding that the planet’s 3D temperature structure has a more profound effect on spectral line shapes than in transmission. In a transmission spectrum, the line strength is determined by the altitude difference between the pressure probed by the line core and the pressure probed by the continuum (e.g., Wardenier et al. 2022, 2023). In an emission spectrum, however, the line strength is directly set by the thermal flux difference111In thermal emission, the flux difference between a line core at wavelength and the adjacent continuum would be proportional to – , with the Planck function and and the temperatures probed by the line core and the continuum, respectively. between these two pressures (e.g., Van Sluijs et al. 2023). If the line core probes a higher temperature than the continuum (due to a thermal inversion), the spectrum contains emission lines. If the temperature profile is non-inverted, the corresponding emission spectrum will contain absorption lines. Because an emission spectrum is the integrated flux over the entire planet disk (which contains different portions of the dayside and nightside at different orbital phases), it will originate from regions with different vertical temperature gradients, giving rise to line shapes that cannot be fully captured by 1D atmospheric models (e.g., Beltz et al. 2021).
The GCM studies also predicted how the Doppler shifts of emission spectra vary in the planet rest frame as a function of orbital phase. In principle, the Doppler shifts of a tidally locked planet should follow a “quasi-sinusoidal” behavior (Zhang et al. 2017; Harada et al. 2021; Malsky et al. 2021; Beltz et al. 2022) as the dayside rotates towards the observer in pre-eclipse (causing a blueshift) and away from the observer in post-eclipse (causing a redshift). Just like in transmission, these anomalous Doppler shifts should naturally lead to peak offsets in – maps obtained from UHJ emission observations. Furthermore, the combination of tidally locked rotation and the 3D temperature structure should cause line strengths to vary substantially along the orbit.
Indeed, numerous emission studies of UHJs have reported species whose cross-correlation signals deviate from the known planetary and/or values, or for which the signal strengths change between different orbital phases (e.g., pre-eclipse vs post-eclipse). This includes observations of Fe, CO, and H2O on WASP-76b (Yan et al., 2023; Costa Silva et al., 2024), CO, H2O, OH, and various metal species on WASP-121b (Smith et al., 2024a; Hoeijmakers et al., 2024; Pelletier et al., 2025; Bazinet et al., 2025), CO, H2O, and OH on WASP-18b (Brogi et al., 2023; Yan et al., 2023), Fe and CO on WASP-189b (Yan et al., 2020, 2022a; Lesjak et al., 2025), Fe, CO, and TiO on WASP-33b (Nugroho et al., 2017; Cont et al., 2021, 2022; Herman et al., 2022; Yan et al., 2022a; Van Sluijs et al., 2023; Finnerty et al., 2023; Mraz et al., 2024), Fe, Fe+ and Cr on KELT-20b (Yan et al., 2022b; Borsa et al., 2022), and Fe on KELT-9b (Pino et al., 2020, 2022).
| Parameter | Value | |
| Orbital period | 1.5637 s (1.81 days) | |
| Pressure range | 200 – 210-6 bar | |
| Radius at bottom | 1.3038 m (1.824 RJup) | |
| Gravity | 7.6 m/s2 | |
| Horizontal resolution | C32 | |
| Vertical resolution | 53 layers | |
| Metallicity and C/O | 1 solar | |
| Drag timescale | ||
| Radiative transfer | non-grey (see Kataria et al. 2013) |
The aim of this work is to build on previous GCM modeling studies and further investigate how the “3D-ness” of UHJs impacts high-resolution observables in thermal emission, such as peak offsets in – maps and phase-dependent signal strengths. We will also assess the detectability of nightside spectra of UHJs. The structure of this manuscript is as follows. Section 2 summarizes our GCMs and 3D radiative-transfer framework. In Section 3, we present and discuss our results. Finally, Section 4 provides a summary and conclusion.
2 Methods
2.1 Global circulation models
In this work, we focus on three 3D atmospheric models of the canonical UHJ WASP-76b ( 2200 K; West et al. 2016), generated with the SPARC/MITgcm (Showman et al. 2009). These are the same models that were used in Wardenier et al. (2021, 2023) to study the phase-dependent behavior of the planet’s transmission spectrum at high resolution. The SPARC/MITgcm is a state-of-the-art, non-grey GCM that has been used to explore many aspects of the atmospheric physics and chemistry of hot gas giants over the last fifteen years (e.g., Fortney et al. 2010; Showman et al. 2013; Kataria et al. 2013; Parmentier et al. 2016, 2018; Steinrueck et al. 2021; Tan et al. 2024; Roth et al. 2024).
Table 1 provides an overview of the most important parameters of our WASP-76b models. For further information about how the GCMs were run, we refer the reader to Parmentier et al. (2018) and Wardenier et al. (2021). In contrast to some previous works that focused on cooler hot Jupiters (Harada et al. 2021; Malsky et al. 2021), we do not consider the effects of clouds. In the context of UHJs, such an approach is justified as their daysides, which emit the bulk of the planet’s thermal flux, are expected to be completely cloud free (e.g., Komacek et al. 2022). The setup of each of the WASP-76b models is identical. The only parameter that is varied is the (uniform) drag timescale , such that we obtain atmospheres with no drag (), weak drag ( s), and strong drag ( s), respectively. The drag timescale represents the typical time it takes for a parcel of air to lose a significant fraction of its kinetic energy. It accounts for a number of physical processes, such as turbulent mixing (Li & Goodman 2010), Lorentz-force braking of ionized winds in the planet’s magnetic field (Perna et al. 2010a), or Ohmic dissipation (Perna et al. 2010b). From an observational perspective, is an important parameter as it governs the effective wind speeds, and thereby the efficiency of heat redistribution, as well as the Doppler shifts observed in the associated spectra. While the drag-free model has a super-rotating equatorial jet (Wardenier et al., 2021), jet formation in the other two models is suppressed, leading to a wind profile that is mainly composed of a day-to-night flow.
Before calculating the thermal emission spectra associated with our WASP-76b models, we first map the GCM outputs onto a 3D grid with altitude (instead of pressure) as a vertical coordinate. The details of this procedure are described in Section 2.1 in Wardenier et al. (2023). Because the dayside has a larger scale height than the nightside due to its higher temperature and lower mean-molecular weight (resulting from the thermal dissociation of H), we extrapolate the nightside down to arbitrarily low pressures to make sure that the atmosphere reaches the same altitude on both hemispheres. The contribution from these pressures to the emitted flux is virtually zero. After binning and extrapolating, our models consist of 326458 105 cells in latitude, longitude, and altitude, respectively.
2.2 Computing emission spectra with gCMCRT
We use the 3D Monte-Carlo radiative transfer code gCMCRT (Lee et al., 2022b) to simulate phase-dependent thermal emission spectra for each of the GCM outputs. As shown in Fig. 1, we calculate spectra at 36 orbital phases , which are separated by 10-degree intervals. Because the planet is tidally locked, different portions of the dayside and nightside are in view at different orbital phases. We do not consider the effects of clouds or multiple scattering in the radiative transfer to avoid overcomplicating the model.
The full inner workings of gCMCRT are described in Lee et al. (2022b), so we will just summarize the most important steps below. For a certain wavelength , each atmospheric cell has a luminosity given by
where is a Doppler-shifted opacity to account for the fact that the emitting cell is moving at a velocity with respect to the observer (due to planet rotation and the wind profile; see Wardenier et al. 2021):
In the above equations, the cell’s mass density, the cell’s volume, is the Planck function, the cell’s temperature, and is the speed of light. The opacity has units cm2/g. Note that the line-of-sight velocity is different for each cell at each orbital phase because of the change in viewing geometry as the planet moves along its orbit222The derivation of the formula for can be found in Appendix B in Wardenier et al. (2021). All spectra are calculated in the planet rest frame, without Doppler shifts due to orbital motion..
In traditional Monte-Carlo radiative transfer, the number of photon packets that are “sourced” from a cell is proportional to the cell’s luminosity, i.e., . However, in the context of UHJs, this approach is not ideal as the large temperature contrast would cause the nightside to be undersampled relative to the dayside, leading to stochastic noise in the modeled spectrum. To circumvent this issue, gCMCRT uses “composite biasing” (Baes et al. 2016; Lee et al. 2017), such that the number of photon packets per cell becomes
Here, is a scaling parameter and is the total number of atmospheric cells in the model. Furthermore, is the total luminosity of the atmosphere and is the total number of emitted photon packets at the given wavelength ( in this work). Equation 3 is a compromise between non-biased Monte-Carlo radiative transfer () and uniform sampling (), where each cell emits the same number of packets. In this work, we use . To ensure that the total luminosity of the atmosphere is conserved, each packet is assigned a weight calculated as
with the average luminosity of all cells. The second equality holds for any value of .
Once a photon packet is sourced from a random position in cell , gCMCRT evaluates the total optical depth encountered by the packet as it escapes from the atmosphere along the line of sight. To account for the effects of winds and planet rotation, we Doppler-shift the opacities in each atmospheric cell traversed by the photon packet according to the local line-of-sight velocity. This calculation involves the same equations as those presented in Section 3.3 in Wardenier et al. (2021). The resulting “peel-off” weight of the photon packet as it exits the atmosphere is (e.g., Lee et al. 2017)
with a factor accounting for isotropic emission. Finally, the flux escaping from the planet disk (with units erg/cm2/s/cm) into the observational direction is obtained by performing a sum over the photon packets emerging from all cells:
with the planet radius corresponding to the top of the atmosphere. To make our computation more efficient, we do not source photon packets from cells that are more than away from the top of the atmosphere along the line of sight. These cells are either situated in very deep layers of the atmosphere or on the “far side” of the planet that is not in view.
2.3 Wavelength range, resolution, and opacities
We simulate the thermal emission spectra of our models across the ESPRESSO (0.38–0.79 m) and IGRINS (1.43–2.42 m) bandpasses. These are the same wavelength ranges that were used by Wardenier et al. (2023) to compute transmission spectra of WASP-76b. The optical spectra are simulated at a spectral resolution = 300,000 ( the ESPRESSO resolution) and the infrared spectra at = 135,000 (3 the IGRINS resolution). Additionally, we compute a few individual spectral lines at = 500,000 to resolve their shapes in more detail (see Figs. 5 to 8).
In both bandpasses, we also use the same opacities as Wardenier et al. (2023). These include continuum opacities due to H2, He, and H scattering, collision-induced absorption (CIA) by H2-H2 and H2-He, and bound-free and free-free transitions of H- (for references, see Table 2 in Lee et al. 2022a), as well as line absorption from Fe, Fe ii, K, Na, Ti, Mn, Mg, Cr, Ca ii (all from Kurucz & Bell 1995), TiO (Mckemmish et al. 2019), VO (McKemmish et al. 2016), CO (Li et al. 2015), H2O (Polyansky et al. 2018), OH (Rothman et al. 2010), CH4 (Hargreaves et al. 2020), CO2 (Rothman et al. 2010), HCN (Barber et al. 2014), and NH3 (Coles et al. 2019). For further details about the opacities and the radiative transfer, we refer to Section 2.2 in Wardenier et al. (2023).
2.4 Cross-correlation and K–V maps
Once we obtained all spectra as a function of orbital phase, we compute the 2D cross-correlation function (CCF; Snellen et al. 2010) by multiplying the spectra with a template shifted by different velocities :
where is the number of wavelength points. To compute the template of a species Fe, CO, H2O, OH, we only include the line absorption from plus continuum opacities in the radiative transfer.
For each species, we generate two templates, which are derived from the drag-free model: the static dayside spectrum at and the static nightside spectrum at . To obtain these “static” spectra, we set the line-of-sight velocities to zero in the radiative transfer (this means that we also remove the effect of planet rotation from the templates). Because the templates do not depend on orbital phase, they do not account for the fact that different parts of the 3D temperature structure are in view at different phases. Hence, the templates are essentially one-dimensional. Before computing the CCF, we subtract the continuum from both the spectra and the templates by convolving them with a Gaussian kernel and subtracting the result. After this, we “standardize” the individual templates by subtracting the mean and dividing by the standard deviation. We evaluate the CCF for km/s, with steps of 0.25 km/s.
For each CCF map, we calculate four – maps associated with the pre-transit (), post-transit (), pre-eclipse (), and post-eclipse (), respectively. To this end, we co-add the CCF values along orbital trails of the form between the specified orbital-phase angles (see also Wardenier et al. 2023). This gives
where SNR is the value of the – map at the point , is a scaling factor such that the maximum is 1, and is the number of emission spectra over which the sum is performed. As shown in Fig. 1, for each of the orbital phases that we consider. At each phase angle, we obtain the CCF value at by interpolating between the two values at the nearest velocity shifts in the CCF map.
3 Results & Discussion
3.1 Global circulation models
Fig. 2 shows the temperature structure and the abundance distributions of the three WASP-76b models. The plots depict the 2D planet disk at four different orbital phases and two different pressures ( and bar). In the IGRINS bandpass, the emission spectrum will typically probe pressures as high as bar (e.g., Fig. 2 in Smith et al. 2024a). In the optical, Fe lines can probe lower pressures in the millibar regime (e.g., Fig. 4 in Pino et al. 2020). In Appendix A, we present individual contribution functions for the chemical species considered in this work. Fig. 3 shows the 1D temperature profiles of the three models in the equatorial plane. As illustrated in both Figs. 2 and 3, the day-night temperature contrast of the atmosphere increases sharply with altitude. At bar, the dayside temperature is still low enough for H2O to persist. However, at bar, the dayside is completely depleted of H2O due to thermal dissociation. On the other hand, the nightside temperature decreases with altitude, resulting in Fe being condensed out of the atmosphere at bar. Owing to its strong triple bond, CO is neither prone to condensation nor dissociation, leading to a uniform 3D abundance (e.g., Beltz et al. 2022; Savel et al. 2022). Consequently, any changes in the line strength of CO should directly probe the planet’s 3D temperature structure.
Finally, Fig. 3 shows that the nightside temperature profile becomes more isothermal with increasing drag strength. This can be understood intuitively. Stronger drag leads to less efficient heat redistribution, and thus less energy being deposited onto the nightside. Furthermore, the radiative timescale (the time it takes for a parcel of air to lose a substantial fraction of its energy by radiation) is proportional to , with the photospheric temperature (Perez-Becker & Showman 2013; Parmentier et al. 2021). Because the nightside photosphere is cooler and radiating less energy into space per unit time, radiative equilibrium will converge to a shallower vertical temperature gradient. As we will demonstrate in the next sections, the strength of the vertical temperature gradient impacts the “detectability” of the nightside spectrum via high-resolution observations.
Fig. 4 shows the line-of-sight velocities in each model due to planet rotation and the 3D wind profile. In the drag-free model (top row), the equatorial jet is clearly visible at both pressures. can reach values of 10 km/s. As the drag strength increases, the wind speeds decrease and the line-of-sight velocities become dominated by solid-body rotation (bottom row). The equatorial rotation velocity of WASP-76b is about 5.3 km/s.
3.2 Individual spectral lines
Figs. 5 to 8 show individual spectral lines of Fe, CO, H2O, and OH as a function of orbital phase in the rest frame of WASP-76b. Each plot depicts the lines associated with the drag-free model (top row) and the strong-drag model (bottom row), respectively. The spectral lines of the static atmospheres (with = 0 km/s) are shown in the background for reference. The first thing to note is that the lines accounting for Doppler effects have a lower line strength compared to their static counterparts. This is a result of Doppler broadening, which is mostly driven by planet rotation. Also, because the line-of-sight velocities in the drag-free model have a larger dispersion due to the wind profile (see Fig. 4), the lines in the spectrum of the drag-free model are slightly broader compared to the strong-drag model. As expected, the emission lines of all species are strongest around the eclipse, when the full dayside of the planet is in view.
Fig. 5 depicts a single Fe line. At every orbital phase, it appears in emission (that is, the flux from the line core is larger than the flux from the continuum), demonstrating that Fe exclusively probes the thermally inverted dayside of the planet. This result is compatible with Fe condensation on the nightside (see Fig. 2). During post-transit and pre-eclipse, the emission lines are blueshifted in the planet rest frame as the dayside rotates towards the observer. During post-eclipse and pre-transit, the dayside rotates away from the observer, inducing a redshift. Remarkably, we can still see weak emission lines at orbital phases , which must originate from a very thin slice of dayside atmosphere that is still in view.
Fig. 6 shows an example of a CO line. Contrary to Fe, the abundance of CO is uniform across the atmosphere, meaning that the CO signal should probe both the dayside and the nightside. During pre- and post-eclipse, the CO lines appear purely in emission, as the strongest signal emerges from the dayside. During pre- and post-transit, however, the spectral-line shapes are more intricate. In the drag-free model, the lines consist of a component in absorption (associated with the nightside) and a component in emission (associated with the dayside). The components exhibit opposite Doppler shifts as the dayside and the nightside have different line-of-sight velocities due to planet rotation. Furthermore, the CO lines in pre- and post-transit appear as each other’s “mirrored” versions, as the dayside is redshifted before the transit, but blueshifted after the transit (and vice versa for the nightside). At phases , the absorption and emission components are roughly equally strong, even though a much larger fraction of the nightside is in view. Closer to the transit, the lines appear mainly in absorption. However, the absorption features are substantially weaker than the emission lines around the eclipse. This is a result of the lower nightside temperature and weaker vertical temperature gradient. A more extreme case is the strong-drag model, in which the absorption features of CO fully disappear. This is because the strong-drag model has a more isothermal temperature profile on its nightside than the drag-free model (see Fig. 3). Therefore, in reality, the detectability of a planet’s nightside absorption features via high-resolution spectroscopy (e.g., Mraz et al. 2024; Yang et al. 2024) will very much depend on the slope of the nightside temperature profile, set by the efficiency of heat redistribution across the atmosphere.
Fig. 7 shows a single H2O line. While H2O is present on both the dayside and the nightside of the planet, it is dissociated at lower pressures on the dayside due to the thermal inversion. Again, we can see that H2O appears in emission during pre- and post-eclipse. As with the CO lines, the strongest differences between the drag-free and the strong-drag model occur during pre- and post-transit. In the drag-free model, the (“V-shaped”) absorption features associated with the nightside are clearly visible. However, in the strong-drag model, the line strength of the absorption component of the H2O lines is very marginal. Even though H2O is abundant on the nightside, the shallow gradient of the temperature profile can wipe out its spectral features as the black-body temperatures probed by the line core and the continuum are very similar. Unlike Fe and CO, H2O does not show up in emission at in the strong-drag model. This is likely an effect of thermal dissociation: H2O lines probe deeper layers on the dayside that are geometrically masked by the nightside at these orbital phases.
Finally, Fig. 8 shows an example of an OH line, whose behavior is very similar to that of the Fe line in Fig. 5. It is worth mentioning that the line strength of the strong-drag model is not only higher due to the smaller effect of Doppler broadening, but also because it has a stronger dayside temperature inversion than the drag-free model (see Fig. 3).
3.3 Phase dependence of the line strengths
In this section, we study the (average) line strengths of different species as derived from their CCFs, rather than focusing on single spectral lines. The line strengths reported in Figs. 9 and 10 are equal to the value of the CCF peak at a given orbital phase. All CCFs were obtained using the same dayside templates, which are described in Section 2.4.
Fig. 9 shows the phase-dependent line strengths of CO and H2O in the drag-free and strong-drag models of WASP-76b, respectively. The left panel shows the line strengths when only accounting for the 3D temperature structure of the atmosphere, but not for Doppler shifts (similar to the “static” lines in Figs. 5 to 8). In the drag-free model, H2O lines appear in absorption during the majority of pre- and post-transit. For CO, the emission features are dominant over a much larger part of the orbit. In the strong-drag model, CO even appears in emission during the entire orbit, which is in agreement with the (grey) spectral lines shown in the bottom row of Fig. 6. These results further underscore the challenges associated with detecting the nightside spectrum of UHJs. Not only are the absorption features constrained to smaller parts of the orbit (if the line strengths become negative at all), they are also weaker than the emission features visible during pre- and post-eclipse. Thirdly, the total thermal flux emanating from the nightside is at most - of the dayside flux, resulting in lower photon counts. All things considered, detecting absorption features associated with UHJ nightsides may well be an order of magnitude harder than detecting emission features from their daysides, especially since many UHJ observations are consistent with atmospheric drag (e.g., Kreidberg et al. 2018; Arcangeli et al. 2019; Coulombe et al. 2023; Wardenier et al. 2024; Demangeon et al. 2024). Therefore, their nightside temperature profiles can be expected to be relatively isothermal.
The right panel of Fig. 9 shows the line strengths of CO and H2O when also accounting for Doppler shifts in the radiative transfer. The main difference compared to the left panel is that there are now certain orbital phases at which the CCF has both a positive and a negative peak (which occur at different effective Doppler shifts due to the rotation of the planet). This coexistence of absorption and emission features can also be seen in the spectral lines of CO and H2O in Figs. 6 and 7.
Another takeaway from Fig. 9 is that the line strengths are not perfectly correlated with the integrated flux from the planet. This is further illustrated in Fig. 10, which shows the line strengths of the four species and the phase curve of the drag-free model in pre- and post-eclipse. To isolate the effect of the 3D temperature structure, we do not include Doppler shifts here. Because the planet has an eastward hotspot offset, the phase-curve maximum occurs before the secondary eclipse. However, the line strengths of CO, H2O, and OH are strongest after the eclipse. The fact that all colored curves are offset from the black phase curve shows that the line strength is not a measure of the absolute temperature probed by the spectrum. Rather, the line strength is set by the vertical temperature gradient of the atmosphere. This result confirms earlier findings by Van Sluijs et al. (2023) based on models and observations of WASP-33b.
3.4 CCF maps of the full orbit
Fig. 11 shows the CCF maps for all GCM outputs of WASP-76b. All maps were obtained by cross-correlating the model spectra with dayside emission templates from the drag-free model. For each of the four species, we recover the “quasi-sinusoidal” dependence of the net Doppler shifts on orbital phase (driven by planet rotation), as identified by Zhang et al. (2017).
All CCFs show very similar behavior between the first and third quadratures of the orbit (i.e. ). Around the first quadrature, all models show a strong blueshift (RV 0) in the planet rest frame. This is because all visible parts of the dayside rotate towards the observer. As a larger part of the dayside comes into view, some regions will become redshifted (see Fig. 4), leading to a smaller net blueshift. As the planet approaches the eclipse, the average Doppler shift converges to zero. After the eclipse, the opposite effect occurs. At the third quadrature, the dayside is exclusively rotating away from the observer, producing a strong redshift (RV 0).
The second column of Fig. 11 shows the CCFs when only accounting for Doppler shifts due to planet rotation. In this case, the net Doppler shifts never exceed the rotational velocity of the planet at the equator, which is indicated by the dashed lines (5.3 km/s for WASP-76b). When the wind profile is taken into account, the Doppler shifts can be stronger. As shown in Fig. 11, the wind profile causes the spectra to be more blueshifted on average. In Appendix A, we plot the phase-dependence of the full width at half maximum (FWHM) for all species and all models in pre- and post-eclipse.
In pre- and post-transit, the behavior of the Doppler shift depends on whether a species also probes the nightside. For Fe and OH, the CCF continues to track the average line-of-sight velocity of the dayside regions that are in view. For CO and H2O, however, the trail “jumps” from the dayside to the nightside once the absorption features in the spectrum (which produce negative CCF values) are stronger than the emission features. We will further explore the nightside signals of CO and H2O in Section 3.7.
3.5 – maps in pre- and post-eclipse
Fig. 12 shows the – maps for the pre- and post-eclipse phases of the drag-free model. The maps were obtained by applying equation 8 to the CCF maps in the left column of Fig. 11. As we consider less than a quarter of the orbit when computing the pre- and post-eclipse maps, we recover the characteristic “degeneracy” between and that is also seen in real datasets (e.g., Birkby et al. 2013; Nugroho et al. 2017; Line et al. 2021; Brogi et al. 2023). At both orbital phases, all species show negative offsets between 3 and 7 km/s. The shifts along the axis are at most 1-2 km/s. When combining the signals from the pre- and post-eclipse (right column in Fig. 12), we observe two changes. Firstly, the – degeneracy disappears, as we are sampling a larger part of the orbit. Secondly, the only significant peak offsets that remain are those along the axis. The value of is negligible in each of the four maps. This is what we expected based on Fig. 11: is the offset of the trail at the eclipse (), which lies close to zero in each scenario. , however, describes the rate of change of the Doppler shifts with orbital phase. Its value is negative because the planet trail goes from being blueshifted to being redshifted in the planet rest frame, counteracting the Doppler shifts due to orbital motion.
Fig. 13 shows the – offsets that we find for the full set of models. Again, all species show a negative offset, driven by planet rotation. Looking at the combined pre- and post-eclipse signals (indicated by the circle markers), we note that the drag-free model produces stronger offsets than the models that include drag. This is because the drag-free model has an equatorial jet that acts in the same direction as the tidally locked solid-body rotation (see Fig. 4). Also, within the drag-free model, the value for Fe (ca. 5 km/s) is larger than the value for H2O (ca. 3 km/s). This is because Fe lines probe lower pressures on the dayside where H2O is thermally dissociated. At these lower pressures, the wind profile is substantially different (top row in Fig. 4), with the dayside being more blueshifted in pre-eclipse and more redshifted in post-eclipse. In the models with drag and rotation only, the values of all species are clustered around 2-3 km/s, which is roughly half the equatorial rotation velocity of WASP-76b. Finally, we note that never exceeds km/s for any species in any of the explored scenarios.
In Fig. 14 we show the ratio between the peak strengths of the – maps in pre- and post-eclipse for all combinations of GCM outputs and species. The left panel conveys the same information as Fig. 10: in the drag-free model, the Fe emission lines are strongest before the eclipse, while the line strengths of CO, H2O and OH peak after the eclipse. However, as the drag strength increases, the dayside temperature structure becomes more symmetric around the substellar point (e.g., Fig. 3), and the ratio between the peak strengths should converge to unity for all species. We recover this behavior when only considering Doppler effects due to planet rotation. Yet, if we also take Doppler effects due to the wind profile into account, the picture becomes somewhat more nuanced. In the drag-free and strong-drag models, the wind profile does not significantly affect the ratio between the signal strengths. However, in the weak-drag model, the wind profile causes the CO and OH lines to become relatively stronger before the eclipse. This implies that the dispersion of the wind speeds probed during post-eclipse must be larger compared to pre-eclipse, such that the emission lines are subject to stronger Doppler broadening, which lowers the line strength after the eclipse. To summarize, Fig. 14 highlights that the 3D temperature structure and the wind profile of a planet can impact the phase-dependent line strengths in opposite ways, warranting caution when it comes to the interpretation of observations.
3.6 Estimating the offset of the dayside emission signal due to planet rotation
Our finding that the planet signal should occur at lower values compared to the “expected” orbital velocity333The Keplerian orbital velocity of a planet on a circular orbit is given by , where is the semi-major axis, is the orbital period, and is the orbital inclination. is in agreement with high-resolution emission observations of different UHJs such as WASP-18b (Brogi et al. 2023), WASP-33b (Cont et al. 2022), and WASP-121b (Hoeijmakers et al. 2024; Bazinet et al. 2025), as well as the JWST/NIRSpec444The maximum spectral resolution of NIRSpec is about 3,000, which is high enough to detect an atmosphere using cross-correlation (see also Esparza-Borges et al. 2023). phase curve of WASP-121b (Sing et al. 2024). When it comes to WASP-76b, Costa Silva et al. (2024) did not find a offset for Fe in pre- and post-eclipse spectra from ESPRESSO. Instead, they reported a blueshift along the axis of 4.70.3 km/s (we will further discuss this observation in Section 3.9). On the other hand, Yan et al. (2023) did report negative offsets for CO and H2O, albeit with large error bars.
Hoeijmakers et al. (2024) detected different metals in the atmosphere of WASP-121b at km/s (note that the equatorial rotation velocity of WASP-121b is about 2 km/s higher than that of WASP-76b). To interpret this observation, they presented an analytical model to calculate Doppler shifts associated with points on the equator (assuming an edge-on orbit, = 0 km/s, and no winds):
In this equation, is the line-of-sight velocity at a point with longitude at orbital phase . It is a combination of the orbital motion of the planet (first term) and its tidally locked rotation (second term). Furthermore, is the orbital period, is the semi-major axis, and is the planet radius.
In the limit where all the flux emerges from the substellar point (), the line-of-sight velocity becomes
where is the Keplerian orbital velocity of the planet, is the rotational velocity at the equator, and is the “observed” based on the emission spectrum. In the above scenario, the offset is equal to the equatorial rotation velocity of the planet. In reality, the signal will not only come from the substellar point. Therefore, for a species that is distributed more or less uniformly across the dayside, should be seen as an upper limit: (see also Fig. 13). When also considering the planet’s wind profile, the inequality can be generalized to , with the equatorial jet speed. We note that these upper limits on are much lower than in transmission, where the offsets along the axis can be multiple times the equatorial rotation velocity of the planet555Equation 8 from Wardenier et al. (2023) yields a estimate of 21 km/s for the transmission spectrum of WASP-76b, which is much larger than the equatorial rotation velocity of 5.3 km/s. (Wardenier et al. 2023).
As demonstrated by Cont et al. (2021), who detected TiO in the emission spectrum of WASP-33b ( 2700 K) at km/s, there exist certain scenarios in which the offset of a species can be positive. When a species is depleted near the hotspot of the planet (e.g., due to thermal dissociation), the strongest signals will emerge from regions near the morning and evening terminators (at ). Consequently, the Doppler shifts pick up an additional redshift/blueshift during pre-/post-eclipse due to planet rotation, leading to a larger apparent orbital velocity (see Figs. 8 and 9 in Cont et al. 2021). In our models of WASP-76b, however, we do not encounter this behavior (also not for TiO; see Fig. 15). This is likely because the equilibrium temperature of WASP-76b is 500 K lower than that of WASP-33b.
3.7 – maps in pre- and post-transit
As demonstrated in Sections 3.2 and 3.3, interpreting the planet signals during pre- and post-transit is more challenging for two main reasons: (i) the signals are much weaker than in pre-/post-eclipse due to the shallower nightside temperature gradient and lower flux levels, and (ii) the spectra contain a mix of absorption and emission features. In theory, one could cross-correlate the planet spectra with phase-dependent templates derived from a 3D model (e.g., Beltz et al. 2021, 2022). In reality, however, the true underlying temperature structure of the target is unknown, so the 3D template will never be a perfect representation of the spectrum. Therefore, to keep things simple and intuitive, we will focus on measuring Doppler shifts using a nightside template (see Section 2.4) that does not depend on orbital phase and only contains absorption lines.
Fig. 16 shows the – maps of CO and H2O for the pre- and post-transit phases of WASP-76b when cross-correlating the spectra with a nightside template. In Fig. 17, we show the CCF maps with the orbital trails corresponding to the – peaks from Fig. 16 plotted on top. The layout of both figures is the same, such that panels in the same rows and columns correspond to each other. Furthermore, we do not discuss the signals of Fe and OH here as these species only probe the dayside of the planet. If detectable during the pre- and post-transit, their emission lines will simply track the dayside regions that are still in view (see Fig. 11).
In the drag-free model, the – maps of CO display both a positive and a negative peak. This is because the underlying CCF map contains both positive and negative values in pre- and post-transit (see first row in Fig. 17). Therefore, when co-adding the CCF values along the phase axis, there exist two well-defined trails along which the sum of the CCFs is either maximized (capturing nightside absorption features) or minimized (capturing dayside emission features). As illustrated in Fig. 17, these trails do not adequately describe the true planet signal (in yellow) over the whole phase range that is considered – they only match the signal at very specific phases. In fact, co-adding the CCF maps along the actual planet trail would lead to a weaker aggregate signal, as the positive and negative CCF values will (partly) cancel each other out. Additionally, we note that the structure of our post-transit – map for CO is reminiscent of Fig. 2 from Mraz et al. (2024), who claim a detection of CO on the nightside of WASP-33b666If robust, these observations would point to little drag and relatively efficient heat redistribution throughout the atmosphere of WASP-33b..
When combining the pre- and post-transit signals of CO, the positive and negative peaks in the – map acquire opposite offsets. This is because the dayside rotates away from the observer in pre-transit and towards the observer in post-transit (), while the opposite holds for the nightside (). As shown in the right column of Fig. 17, we need an orbital trail with a positive amplitude (in red dashes) to maximize the sum of CCF values, while we need a trail with a negative amplitude (in blue dashes) to minimize this sum. However, since the trails are only set by two parameters, neither of them is able to fully reproduce the more intricate phase dependence of the true planet trail.
In the strong-drag model (second row in Figs. 16 and 17) the – map of CO only features a negative peak, indicating that dayside emission features are dominant across the whole pre- and post-transit phase. This behavior is in agreement with the phase dependence of the line strengths plotted in Fig. 9. In this scenario, we are thus only probing the dayside of WASP-76b.
In contrast to the CO signal, the H2O lines in the drag-free model are visible in absorption during the entire pre- and post-transit (third row in Figs. 16 and 17). As a result, the – map only contains a positive peak. Furthermore, because the (yellow) planet trail does not suffer from discontinuities or sign changes of the CCF values, the and values associated with the peak are a very good description of the actual planet signal.
In the strong-drag model (bottom row in Figs. 16 and 17), the behavior of the H2O signal is very similar to that of the CO signal in the drag-free model. Because the nightside of this model is more isothermal, H2O absorption is limited to a narrower range of phases. Hence, the – maps feature two peaks again, which are associated with the dayside and the nightside, respectively.
A final aspect to highlight is that the negative – peaks (associated with dayside emission) in the individual pre- and post-transit maps tend to have strong offsets. This is because, mathematically, is the radial velocity at which the corresponding trail intersects the line in the CCF map. The negative peaks, however, encode information from the dayside, whose signal gets stronger at phases further away from . Therefore, the value should not be interpreted as an effective line-of-sight wind speed. For example, in the post-transit map of H2O in the strong-drag model, the Doppler shift of the dayside is about 5 km/s. However, the trail that corresponds to the negative peak in the – map intersects the line at km/h.
3.8 A note on measuring offsets from high-resolution observations
To infer the offset of a species from a – map, we need a reference value for the true (“expected”) orbital velocity of the planet. So far, we have assumed that our knowledge of this value is perfect. However, as shown in Fig. 18, the uncertainties in the literature values of are typically multiple km/s for UHJs. As discussed in Section 3.6, we expect the offset of a species in the dayside spectra to be of similar magnitude. Therefore, it will be hard for certain planets to meaningfully measure offsets in emission, unless the errors in their true orbital velocities are reduced.
In general, the Keplerian orbital velocity of a planet should obey the following equation (Torres et al. 2010; Sing et al. 2024):
where is the stellar mass, is the orbital inclination, is a constant, is the orbital eccentricity, is the stellar RV semi-amplitude, and is the orbital period. To lower the uncertainty in , we thus need to reduce the error in the other parameters in this equation.
For UHJs, the eccentricity, inclination angle, and orbital period tend to be very well constrained (e.g., Ehrenreich et al. 2020; Borsa et al. 2021; Wong et al. 2021). Based on dynamical arguments (Rasio et al. 1996), it is often assumed that . So far, observations have been consistent with this assumption within tight error margins. Furthermore, the errors in the orbital period of UHJs are typically less than a second. For transiting planets, the orbital inclination can be constrained from the Rossiter-McLaughlin effect.
This leaves us with and . For a single star, the way to infer its mass is through evolutionary tracks and isochrones (e.g., Schaller et al. 1992; Torres et al. 2010; Sing et al. 2024). Hence, the error in is impacted by model uncertainties, as well as measurement uncertainties of the star’s effective temperature, surface gravity, and metallicity. For WASP-76 and WASP-121, for example, the error bar on the stellar mass is about 0.02 (Tabernero et al. 2021; Borsa et al. 2021).
The value of can be constrained from radial-velocity measurements. The main reason why the uncertainty for WASP-76b is so low is the small error bar on the value of its host star ( m/s; Ehrenreich et al. 2020). For stars such as WASP-121 ( m/s; Bourrier et al. 2020) and TOI-2109 ( m/s; Wong et al. 2021), the error bars are significantly larger, resulting in a higher uncertainty in the expected planetary orbital velocity.
The relatively large uncertainties in the values of some UHJs could be a motivation for more in-depth characterization of their host stars to better constrain their RV semi-amplitude. High-precision spectrographs such as ESPRESSO, MAROON-X, or NIRPS would be ideal for this task. We note, however, that A-type stars such as WASP-33 and WASP-189 are fast-rotating and/or active, which could make it significantly harder to lower the error to the km/s level.
Alternatively, a way to circumvent the observational uncertainties in could be to report differences between the values of two species (say, Fe and H2O), thereby “subtracting out” the planet’s orbital velocity. However, as shown Fig. 13, differences between the values of two species will typically be small (at most 1-2 km/s in the drag-free model), so these measurements would require high precision to be meaningful.
3.9 A note on offsets
Our WASP-76b models do not predict any large offsets in the planet’s dayside emission spectrum, especially when pre- and post-eclipse observations are combined. The reason for this is that the CCF trails in Fig. 11 intersect the line at RV 0 km/s, which is best fit by an orbital trail with km/s. In light of our models, the consistent blueshifts observed in the Fe lines of WASP-76b during pre- and post-eclipse with ESPRESSO ( km/s; Costa Silva et al. 2024) are a bit puzzling777Just before submitting the revised version of this manuscript, we learned that a re-analysis of the ESPRESSO data by another team revealed no shifts in , but rather a negative 5 km/s shift in . This result is much more in agreement with our models (G. Guilly, private communication).. Costa Silva et al. (2024) suggested that this blueshift could be caused by a strong updraft of material around the substellar point of the planet (see their Fig. 8), which effectively moves Fe towards the observer. Our GCMs, however, do not provide strong evidence for such vertical transport, perhaps due to missing physics888One mechanism that our WASP-76b models do not consider is heat transport due to hydrogen dissociation/recombination (e.g., Bell & Cowan 2018; Komacek & Tan 2018; Tan & Komacek 2019; Roth et al. 2021). However, post-processing of the WASP-121b models from Wardenier et al. (2024) (which include hydrogen dissociation/recombination) shows that this additional physics does not lead to stronger offsets in dayside emission spectra. We refer to Appendix B in Bazinet et al. (2025) for the – maps of these models.. Additional forward modeling and follow-up observations with different instruments will be required to resolve the discrepancies between the current data and our models.
One aspect of the observations from Costa Silva et al. (2024) that is more in agreement with our models is the decreasing blueshift of Fe over the course of the orbit: RV = 6.0 0.4 km/s in pre-eclipse, while RV = 3.3 0.5 km/s in post-eclipse. This could be a signature of planet rotation. In principle, such a change in Doppler shift with phase should translate to a (negative) offset in the – map. However, Costa Silva et al. (2024) reported that the Fe emission signal of WASP-76b is consistent with km/s. One reason for this could be differences in the signal-to-noise ratio between the four ESPRESSO visits, such that not all orbital phases contribute equally to the “– fit”.
While Costa Silva et al. (2024) found a negative offset for Fe on WASP-76b, Lesjak et al. (2025) reported positive offsets for Fe and CO on WASP-189b, based on pre- and post-eclipse observations from CRIRES+ ( km/s). Lesjak et al. (2025) attributed their measurements to strong day-to-night winds that impart a redshift on the emission spectrum. With regard to WASP-189b, we note that its equatorial rotation velocity is about 2 km/s lower than that of WASP-76b (see Fig. 18), which leads to smaller offsets due to planet rotation. In this case, it is likely that the wind profile is the dominant driver of the line-of-sight velocities. Along this line of reasoning, one could speculate that UHJs with high rotation rates (i.e. short orbital periods) are more likely to show offsets, while UHJs with lower rotation rates are likely to show stronger offsets. Such trends could be identified by modeling emission spectra for a sample of UHJs.
4 Summary and conclusion
In this work, we simulated phase-dependent emission spectra of the canonical ultra-hot Jupiter (UHJ) WASP-76b, based on 3D atmospheric models from the SPARC/MITgcm. We post-processed the GCMs with gCMCRT to obtain emission spectra at 36 phases along the orbit. We then cross-correlated the spectra with different templates to obtain CCF and – maps for Fe, CO, H2O, and OH. This allowed us to identify how their Doppler shifts and line strengths evolve throughout the orbit. Because of the stark day-night temperature contrasts that prevail on UHJs, the emission spectra undergo drastic changes as different parts of the 3D temperature structure rotate into and out of view.
As the number of high-resolution observations of UHJs is steadily increasing, 3D forward-modeling studies are crucial to fully leverage the information content of current and future telescope data. In a few years’ time, UHJs will undoubtedly be prime targets for the E-ELT (e.g., Palle et al. 2023), which will deliver transmission and emission spectra at unprecedented signal-to-noise ratios. These will allow us to characterize the climate and composition of UHJs in much finer detail. Finally, we note that cross-correlation analyses of UHJs are now also possible with JWST/NIRSpec (Sing et al. 2024), albeit at more modest spectral resolutions (). This means that our work will also be relevant for understanding certain aspects of space-based observations.
We summarize our most important findings below:
Detecting the nightsides of UHJs (through CO and H2O absorption lines) could be an order of magnitude harder than detecting their daysides. This is due to the more isothermal temperature profile of the nigthside, which causes the absorption features to be much weaker than the emission features associated with the (thermally inverted) dayside. In our models with drag, parts of the pre- and post-transit phases are dominated by dayside emission, even though a much larger portion of the nightside is in view. In general, the detectability of the nightside spectrum will depend on the efficiency of heat redistribution across the atmosphere of an UHJ, which sets the vertical temperature gradient on the nightside.
Line strengths in the thermal emission spectrum of a planet are set by the vertical temperature gradient and not by the total flux. For example, our drag-free model of WASP-76b has an eastward hotspot offset (such that the total flux peaks before the eclipse), while the line strengths of CO, H2O, and OH are highest after the eclipse.
In each of our WASP-76b models, the phase-dependent Doppler shifts show a “quasi-sinusoidal” behavior in the planet rest frame that is driven by planet rotation. During pre- and post eclipse, this results in a negative offset in the – map of a species. The upper limit on this offset is given by , with the equatorial rotation velocity of the planet and the equatorial jet speed (in the case of atmospheric drag the last term can be ignored). When combining the pre- and post-eclipse phases, we find negligible offsets (at most 1 km/s) for our WASP-76b models.
In pre- and post-transit, CO and H2O can appear in emission and absorption. If this is the case, the corresponding – maps will feature both a positive and a negative peak. When combining the pre- and post-transit phases, the peaks will be located at opposite offsets that are apart. In the individual – maps of the pre- and post-transit, the peaks can have have larger offsets that are not necessarily a measure for the line-of-sight velocities in the atmosphere.
For many UHJs the uncertainty in the true Keplerian orbital velocity is roughly as big as the offset which their atmospheres could imprint on the emission spectra. If this is the case, it will be hard to meaningfully constrain values and make inferences about the dynamics of the planet. More in-depth RV characterization of UHJ host stars would be beneficial in this regard.
J.P.W. acknowledges support from the Trottier Family Foundation via the Trottier Postdoctoral Fellowship, held at the Institute for Research on Exoplanets (IREx). We also acknowledge support from the Canadian Space Agency (CSA) under grant 24JWGO3A-03. V.P. is funded by the French National Research Agency (ANR) project EXOWINDS (ANR-23-CE31-0001-01). We thank the members of the NIRPS science team for insightful discussions.
Appendix A Supplementary figures
In this appendix, we present two supplementary figures. Figure 19 shows the rough pressures probed by the emission lines of Fe, CO, H2O, and OH, as derived from the 1D hemispheric dayside average of the drag-free GCM of WASP-76b (see the figure caption for further information about methods). The plots demonstrate that different species probe the atmosphere at a range of pressures between 10-1 and 10-5 bar. While H2O probes relatively deep layers of the atmosphere (pressures higher than 10-3 bar), the Fe emission lines mostly emerge from pressures between 10-3 and 10-5 bar.
Figure 20 shows the full width at half maximum (FWHM) of the CCFs from Fig. 11. During pre-eclipse (), the FWHM increases for all species as a larger part of the dayside rotates into view. That is, the emission signal emerges from an increasingly larger region on the planet disk, probing a wider range of line-of-sight velocities. The opposite effect occurs during post-eclipse (), when parts of the dayside rotate out of view. Comparing the different GCM outputs of WASP-76b, we note that the drag-free model produces the broadest CCF peaks. This is because its strong winds, including the equatorial jet, increase the dispersion in the probed line-of-sight velocities (see Fig. 4). Interestingly, at , the FWHMs of the weak-drag model are 2 km/s lower than the FWHMs of the strong-drag model, suggesting that the day-to-night flow partially counteracts the line-of-sight velocities due to planet rotation. For CO and Fe, the weak-drag model also predicts the CCF peaks to be somewhat broader in post-eclipse than in pre-eclipse.